whitney has 20 candies and eats 2 each minute. how many candies will she have in 10 minutes?

Answer by wilft1(217)

we can write this out as an equation...
she starts off with 20 candies, she eats 2 each minute times 10 minutes
sooo...
20 - 2 x 10 =
20 - 20 = 0